Zebra Company reports the following figures for the years ending December​ 31, 2019 and​ 2018:

2019

2018

Net Sales

$64,000

$48,000

Cost of Goods Sold

$41,000

$33,000

Gross Profit

$23,000

$15,000

What are the percentage changes from 2018 to 2019 for Net​ Sales, Cost of Goods Sold and Gross​ Profit, respectively?​ (Round your final answers to one decimal​ place, X.X%)
Answer #1

Percentage change

Net sales = (64,000-48,000)/48,000 = 33.33%

Cost of Goods Sold = (41,000-33,000)/33,000 = 24.2%

Gross profit = (23,000-15,000)/15,000 = 53.3%

33.3%, 24.2%, 53.3%